OpenStudy (anonymous):

isolate the a d=v*t+1/2*a*t^2

OpenStudy (anonymous):

step 1. subtract vt on both sides step 2. multiply both sides by 2 to get rid of 1/2 step 3. divide by \[t^2\]

OpenStudy (anonymous):

vt is being multiplied can we still subtract it? how about we just divide the other side by t and then subtract v1

OpenStudy (anonymous):

here's what i got: 2d-v1/t/t^2 = a

OpenStudy (anonymous):

it's another method of doing it dividing by t leaving you v+1/2*a*t then you subtract v and divide t multiply by 2.

OpenStudy (anonymous):

however, in your case the math you did was wrong due to dividing t and dividing t^2 again leaving the denominator with t^3 where we do not even have t^3 to begin with

OpenStudy (anonymous):

a=2d-v/t^2 is the answer that i got.

OpenStudy (anonymous):

let me confirm

OpenStudy (anonymous):

you forgot to multiply t. or you forgot to put them seperate after you canceled one t

OpenStudy (anonymous):

the correct answer should be a= (2d-2vt)/t^2 or a= 2d/t^2-(2v/t)

OpenStudy (anonymous):

2 is only multiplied with the numerator right? and if so do we only multiply d and v or d,v and t?

OpenStudy (anonymous):

(2(d-vt))/(t^2) = a